Three charged particles are placed at the corners of an equilateral triangle of side 1.2



m. The charges are +4 μC, -8 μC, and -6 μC. Calculate the magnitude and direction of



the net force on each due to the other two.



1
Expert's answer
2022-05-31T10:04:46-0400

"F_1=\frac{kq_3q_2}{r^2}=+300~mN,"

"F_2=\frac{kq_3q_1}{r^2}=-150~mN,"

"F_3=\frac{kq_1q_2}{r^2}=-200~mN."